### v1

# Hanbury Brown-Twiss interference with massively parallel spectral multiplexing for broadband light

Abstract
Two-photon interference in the spectral domain is a powerful resource for quantum technologies, enabling both precision measurements and scalable entanglement distribution. Here, we report the first demonstration of massively parallel, wavelength-resolved photon bunching, revealing Hanbury Brown–Twiss correlations across 100 independent spectral channels. These observations are enabled by a fast, data-driven single-photon spectrometer that achieves 40 pm spectral and 40 ps temporal resolution over a 10 nm bandwidth, allowing simultaneous access to spectro-temporal photon correlations without the need for narrowband filtering. This approach preserves photon flux while enabling high-dimensional quantum interference measurements across a broad spectrum. Our results establish frequency-multiplexed two-photon interference as a scalable and throughput-efficient platform for quantum-enhanced photonic technologies, with applications ranging from stellar intensity interferometry to broadband entanglement swapping in quantum networks.
1 Introduction
Two-photon correlation analysis stands as a cornerstone of contemporary quantum optics, tracing its origins back to foundational studies by Glauber [1, 2]. This powerful method has found diverse applications across numerous fields, including super-resolution microscopy [3, 4], quantum communication protocols [5], two-photon resonance fluorescence [6, 7], stellar intensity interferometry (SII) [8, 9, 10, 11, 12], and its advanced phase-sensitive variants [13, 14, 15, 16]. Central to these applications are distinct correlation signatures between photons arising from fundamental quantum interference effects, predominantly characterized by the indistinguishability of photons as in the Hanbury Brown-Twiss (HBT) [17, 18] and Hong-Ou-Mandel (HOM) [19] phenomena. In this study, we directly measure and quantify photon indistinguishability by employing a fast spectrometer combined with precise spectral and temporal data analysis conducted during post-processing as schematically presented in Figure 1. For the first time, we unambiguously observe the HBT effect of photon bunching in broadband light simultaneously across 100 spectral bins in a continuous 10 nm range, establishing a new methodology for massively multiplexed correlation sensing over an extensive frequency range, with the potential of substantially improving the resolution, sensitivity, and scalability for quantum optical measurements.
Figure 1: Conceptual illustration of wavelength-resolved two-photon interference. Broadband thermal light is split into two beams and analyzed using a dual-arm spectrometer, producing time- and wavelength-resolved single-photon spectra for each path. A spectro-temporal correlation analysis is performed between spectral channels on both sides. When photons occupy the same spectral bin $(\lambda_{i}=\lambda_{j})$, quantum bunching leads to an HBT peak — an enhancement in coincidence counts at small time delays between the two photons. In contrast, mismatched spectral bins $(\lambda_{i}\neq\lambda_{j})$ yield flat, uncorrelated distributions, as expected for distinguishable photons. This setup enables massively parallel observation of two-photon interference across a wide spectral range.
A particularly compelling application of frequency-multiplexed two-photon techniques aims at achieving superior angular (astrometric) resolution in astronomy. Traditionally, interferometric methods have relied on single-photon interferometry, which requires precise photon phase measurements of light emitted by astronomical sources detected by telescope stations positioned along extensive baselines. However, achieving large baselines with single-photon interferometry is technically challenging due to the stringent requirement for phase-stable optical links [20, 21]. In contrast, techniques based on the two-photon HBT effect, historically employed in SII to measure stellar diameters, inherently bypass this requirement. Recent proposals suggest further enhancements using quantum-mechanically entangled photon pairs, which could eliminate the need for phase-stable links altogether, thus enabling arbitrarily long baselines [13, 22, 23, 14, 24]. This quantum-enhanced approach holds transformative potential for astronomical observations by dramatically improving achievable resolutions. Moreover, frequency multiplexing can amplify these benefits, facilitating simultaneous and independent measurements across multiple spectral bins, subsequently combined to enhance measurement sensitivity significantly. Notably, this multiplexed methodology is equally advantageous for conventional SII techniques.
Another application for the spectrometer with excellent spectral and temporal resolutions is resonance fluorescence, where precise two-photon interference and correlation measurements can yield groundbreaking insights into quantum optical processes in atomic and molecular physics. Spectral correlation patterns observed in resonance fluorescence typically manifest as distinct photon bunching lines and anti-bunching circles, unveiling quantum behaviors at off-resonant frequencies [25, 6].
Wavelength-multiplexed two-photon interference is a critical resource for boosting the throughput of quantum communication networks. In widely used entanglement swapping protocols, two spontaneous parametric down-conversion (SPDC) sources each generate a pair of entangled photons. One photon from each pair is directed to a central station where a HOM interference measurement projects entanglement onto the two distant nodes [26]. These protocols are foundational for linking quantum processors and for enabling device-independent quantum key distribution (QKD). However, the probabilistic nature of SPDC imposes a fundamental trade-off between brightness and fidelity: increasing the pair production rate raises the likelihood of double-pair emissions, which degrade entanglement quality. Additionally, SPDC sources typically emit broadband spectra, requiring narrow spectral filtering to ensure photon indistinguishability and high interference visibility [27]. Such filtering, however, severely limits the usable photon flux and network throughput. To overcome this, we propose treating a broadband SPDC source as a collection of parallel, independent wavelength channels — each producing low-probability, Fourier-limited single-photon pairs. By leveraging wavelength multiplexing instead of spectral filtering, this approach can substantially increase the number of successful entanglement swaps per unit time. This parallelization strategy represents a promising route toward scalable, high-rate, all-photonic quantum repeaters [28].
Previous attempts to measure the frequency-multiplexed HBT effect were constrained by either limited temporal or spectral resolution or by a narrow frequency range with few spectral bins. Earlier investigations of spectrally resolved HOM interference employed slower cameras, utilizing nanosecond-scale resolution to identify temporal coincidences in photon pairs. Temporal coherence was usually maintained by precisely delaying one photon relative to the other using mechanical stages, effectively enabling sub-picosecond precision and thereby providing good visibility for HOM interference [29, 30, 31]. More recent efforts, such as those described in [32, 33, 34], achieved significantly improved timing resolution, 20–40 ps. However, these studies still faced limitations, with only a small number of spectral channels, seven channels in [32] and five narrow neon spectral lines in [34], or restricted spectral coverage, less than 0.3 nm, when investigating two-photon transitions in polariton condensates [33]. Similarly, SII experiments typically implemented frequency binning through multiple narrow-band filters and standalone photon detectors, thereby limiting scalability [35, 36, 37, 11].
Other existing techniques, which could achieve simultaneously good spectral and temporal resolutions for the frequency multiplexing employ dispersive fibers to translate changes in time delay to the photon frequency [38], spectral phase shaping using spectral shear via electro-optic modulation [39], and an approach based on chirped fiber Bragg gratings [40]. These systems are interesting in their own right, but typically require complex infrastructure for pulsed operation and synchronization, and observation of the two-photon interference was not their primary motivation.
Maximizing interference contrast, or equivalently, visibility, necessitates simultaneously achieving exceptional temporal ($\sim$10 ps) and spectral ($\sim$10 pm) resolutions, thereby approaching the fundamental limits set by the Heisenberg Uncertainty Principle [41], also known as the Gabor or Fourier or Heisenberg-Gabor limits [42]. Current advances in multi-pixel single-photon detector arrays enable building of wideband spectrometers and also address challenges related to temporal resolution, which commonly limits the achievable contrast due to peak smearing in HBT measurements. These innovative data-driven detectors significantly enhance the practical applicability of quantum optical phenomena across various real-world scenarios, opening numerous opportunities for cutting-edge applications.
In the following, we describe the results demonstrating the HBT effect of two-photon interference for broadband light, post-selecting photons with the same frequency and timestamps as determined by precise measurement in the spectrometer. We then discuss the relevance of the results to several scenarios of two-photon interference with performance projections based on the achieved performance.
2 Methods
We describe below a fast spectrometer with spectral and temporal resolutions of 40 pm and 40 ps, respectively, together with light sources used for the measurements and calibrations. Figure 2 shows the layout of the spectrometer and two light sources used in the measurements. Similar spectrometers have already been described in our previous publications [43, 44, 34], so only a brief recap of their design and essential components follows. One of the main outcomes of this work is that we observed the photon bunching owing to the HBT effect for pairs of thermal photons from a broadband source that were matched in time and in frequency. The photons without frequency matching do not show this behavior.
Figure 2: Layout of dual spectrometer with two light sources: neon lamp and light-emitting diode (LED). For both sources, the light is polarized with a linear polarizer (LP) and filtered through a spectral bandpass filter (SBF) before entering a fiber port (FP), where it is coupled into a single-mode fiber. The source switch ($S$) is essentially a fiber-to-fiber connecting sleeve, which allows for the selection of one of the two sources at a time. A 1-to-2 50:50 fiber-coupled beamsplitter (FBS) splits and sends the light into two arms of the spectrometer. One of the arms is delayed via an additional single-mode fiber (SMF). Both arms are connected to adjustable collimators (AC) that are used for focusing the beams of light. With the help of a mirror (M), the output beams are guided onto a ruled reflective diffraction grating (RRDG). The resulting spectrum is focused via an achromatic doublet (AD) onto the LinoSPAD2 sensor (LS2). Everything besides the two light sources and fiber-coupling components is located inside a light-tight enclosure. All fibers used in this setup are single-mode ones, including the beamsplitter.
2.1 LinoSPAD2 detector
The key part of our spectrometer is the LinoSPAD2 camera. In the LinoSPAD2 sensor [45, 46], each pixel is a single-photon sensitive avalanche photodiode (SPAD) with a $26.2\times 26.2$ µm2 size and 25.1% fill factor. The sensor consists of a linear $512\times 1$ pixel array with a median dark count rate of $\sim 125$ cps/pixel, cross-talk of 0.2% [47, 48] and final photon detection efficiency of about 30% at 520 nm for the sensor equipped with microlenses [45, 49] while operating at room temperature. After calibrations, the temporal resolution was determined to be equal to 40 ps (rms) employing an SPDC source of photon pairs [50]. In this work, we employed 256 pixels (one half of the sensor). LinoSPAD2 has already demonstrated good performance in detecting the HBT photon bunching signatures in our prior experiments with neon and argon spectral lines [44, 34, 47].
2.2 Spectrometer
The light entering the spectrometer is polarized and filtered via a $(640\pm 5)$ nm bandpass filter before entering a fiber port. A 1-to-2 50:50 fiber-coupled beamsplitter splits the light into two beams, directing it to the two arms of the dual spectrometer. The outputs of the beamsplitter are connected to adjustable collimators that help with focusing and improving the spectral resolution. Each beam is then guided onto a ruled reflective diffraction grating with 1200 grooves/mm after reflection in a silver-coated mirror. The diffracted linear image of the spectrum within each arm of the dual spectrometer is focused through a 200 mm focal length lens onto the sensor. Every fiber used in this setup, including the fiber-coupled beamsplitter, is single-mode, and everything except the light sources is located inside a light-tight enclosure. To separate a possible cross-talk from the HBT effect, one of the two signal photons is delayed via an additional 1 m fiber in one of the dual spectrometer arms. This results in a shift of $\sim 5$ ns of the HBT peak relative to $\Delta t=0$ in the coincidence histograms [47, 48].
2.3 Light sources and spectral calibrations
For the measurements presented here, we used two types of thermal light sources: broadband LED and neon lamp, both shown in Figure 2. Figure 3 shows two mirrored instances of a 10 nm wide spectrum in the dual spectrometer. The spectrometer optics were aligned such that the two spectra appeared in the same sensor without overlap. This is directly exploiting the data-driven operation of the LinoSPAD2 single-photon detector, which can independently register multiple photons arriving at the same time with excellent temporal resolution.
Figure 3: Two 10 nm wide spectra in the dual spectrometer. The light source is an LED light passing through a $(640\pm 5)$ nm bandpass filter. Two neon lines, 638.3 and 640.2 nm, measured in the same conditions as the broadband light, are overlaid with the broadband spectra, providing precise calibration of the spectrometers. Several weaker lines are also visible.
The spectrometer was characterized using the neon emission spectrum [51], which has a large number of narrow lines. Figure 3 shows two neon spectra in the dual spectrometer. Two bright neon lines, 638.3 and 640.2 nm, are prominent in both spectrometers. Note that the wavelength is increasing in opposite directions for the two spectrometers due to the orientation of their optical elements. The two neon lines have well-defined wavelengths and can be used to calibrate the correspondence of the wavelength to position in the sensor. The spectral resolution (rms) and pixel scale of the spectrometer was determined to be equal to 0.04 nm and 0.11 nm/pixel, respectively [44].
2.4 Data processing pipeline
Acquisition of the dataset used for the analysis took nine hours, with a fraction of the detector live time with open shutter of about 10%. On average, the photon rate in a pixel with maximum occupancy was 570 kHz, while the total photon rate from all 256 pixels was approximately 71 MHz, corresponding to a data rate of $\sim 270$ MB/s and resulting in 2 TB of raw data.
In the data analysis, pairs of pixels from the two different spectrometers were selected, and their timestamps were compared calculating delays between detections of pairs of photons. These timestamp differences are then used to build a histogram of coincidences, where a peak — or enhancement in coincidences — may correspond to either the cross-talk or HBT effects [50]. The preprocessed timestamp differences were saved in independent datasets to separate this part of the analysis from its final phase involving the fitting. This resulted in another 2 TB of preprocessed data.
3 Results
3.1 HBT peaks
In the data analysis, we select two pixels corresponding to two spectral bins from different spectrometers and find hits in those spectral bins with close timestamps. The HBT effect manifests itself as the enhancement of random time coincidences of thermal photons at a certain time difference, producing a so-called HBT peak. Theoretically, the maximum contrast of the peak, defined as the ratio of the peak height to the average background rate, can be equal to 100%. Figure 4 shows examples of HBT peaks after combining information from two pixels corresponding to two spectral bins (each from one of the spectrometer arms) with the same wavelength in our experiments.
Figure 4: a–d) Time difference distributions in two spectral bins in the dual spectrometer, four combinations in total, with examples of different contrast values for HBT peaks determined by the fit. e) Sum of all diagonal combinations where the two frequencies are expected to be equal. The peak contrast as determined by a fitting procedure is equal to $(2.0\pm 0.1)$% and sigma 70 ps; f) HBT peak after filtering with $(656.40\pm 0.05)$ nm narrow band filter.
To estimate the interference contrast in an unbiased manner, we developed a simple fitting procedure with a minimal number of free parameters and applied it uniformly to all combinations of pixel pairs. The HBT peak was fitted with a Gaussian function with its position predicted by the offset calibration [48] and constrained to $\pm 200$ ps of its nominal value. The Gauss sigma was constrained to the values from 60 to 80 ps, as its possible interval was determined in independent high-statistics measurements. The normalized background was set equal to one. The height of the peak — or contrast — was constrained to the values from -10 to 10%.
Typical cases presented in Figure 4 have the peak contrast values from 1.9 to 3.8%. Figure 4 e) shows the sum of time difference distributions where all bin pairs of equal frequencies with expected non-zero signal are combined. The resulting distribution has a prominent HBT peak with $(2.0\pm 0.1)$% contrast and 70 ps sigma. In this case, non-ideal temporal offset calibrations would decrease the contrast and increase the width due to additional smearing, compared to the fits of individual pairs in Figure 4 a–d). We also note that in this case, we acquired enough statistics to observe that the background is not flat but has a slowly changing shape, which was accounted for by a sine term in the fit for this particular case. Figure 4 d) also shows the HBT peak for the LED broadband light after an ultra-narrowband $(656.40\pm 0.05)$ nm filter was used without a spectrometer in a different set of measurements. Note that the width of the filter approximately matches the bin width in the spectrometer, resulting in similar contrasts. The measured contrast, in this case, is $(3.9\pm 0.1)$%.
3.2 Matrix of contrasts
Finally, Figure 5 shows the full matrix of contrasts for all combinations of pairs of spectral bins in a matrix of $100\times 100$ pixels, so 10,000 combinations in total. We expect HBT peaks to appear only for the cases of equal frequencies for combinations of spectral bins which belong to the diagonal of this matrix as is indeed apparent here. The contrasts were estimated by fitting the time difference distribution as described in detail in the previous subsection. In the case of the off-diagonal elements, where the HBT peaks were not expected, the contrast was close to zero, while for the diagonal elements with HBT peaks, the contrast is about 3 to 4%. We consider this result a remarkable confirmation of expected behavior for the HBT effect massively multiplexed in frequency, and also a sound quality proof of our instrument.
Figure 5: Summary 2D histogram of contrasts extracted by the fitter from histograms of coincidences for each pair of spectral bins. The 1D plots above and on the right show the two spectra in the dual spectrometer. The diagonal corresponding to the spectral bins matched in frequency clearly shows the non-zero contrasts of the observed HBT peaks. As the edges of the two spectra have very low photon rates, this translates to large statistical fluctuations near the 2D histogram edges. The color scale for contrast values was limited from 1-1% to 5% to remove a small number of outliers on the edges. Empty vertical and horizontal lines correspond to several noisy, and therefore masked pixels, and also to dead pixels.
Figure 6 shows contrasts for two different diagonals: the main one corresponding to the matching spectral bins showing contrasts of the HBT peaks and one shifted 2 spectral bins away, corresponding to mismatching bins. A mismatch of just $\sim 0.2$ nm in the wavelength of the two photons leads to a loss of the photon bunching.
Figure 6: Contrast values along two selected diagonals: the HBT diagonal corresponding to equal wavelengths and a diagonal offset by 2 spectral bins (mismatched wavelengths).
4 Discussion
In the following section, we discuss practical implications of our results for astrometric resolution and quantum entanglement-swapping protocols, highlighting pathways for further improvements through advanced detector technologies.
Astrometric resolution: The presented results are an important step forward for the SII and quantum-assisted schemes of optical interferometers [52, 24] because the spectrometer can be directly employed for spectral binning in the corresponding experiments. In the case of the tested spectrometer, we can combine about 200 pairs of spectral bins where the contrast is nonzero, corresponding to three adjacent diagonal lines. The uncertainty of the opening angle between two stars $\Delta\theta$ can be expressed as in [14], showing dependence on the most important variables:
$\sigma\left[\Delta\theta\right]\propto\frac{\lambda}{B}\,\frac{1}{V}\,\frac{1}{T}\frac{1}{\sqrt{\bar{n}T}}.$ (4.1)
It is proportional to the wavelength $\lambda$ and inversely proportional to the baseline between the interferometer stations $B$, allowing us to gain from longer baselines. The temporal and spectral resolutions of the spectrometer are accounted for by the visibility $V$, while other photon collection and detection parameters, such as the size of the mirrors and the detection efficiency, are accounted for in the detected rate of photons $\bar{n}$. With $T$ being the length of observation period we also note that the overall $T^{-3/2}$ dependence is more favorable than simply the expected $T^{-1/2}$ gain from photon pair statistics, reflecting the advantage of being able to use the measurement of a fringe oscillation rate due to the Earth rotation as proposed in [14].
The contrasts presented in Figure 5 have direct correspondence to $V$ and therefore in the presented study, we directly measured the visibility for each pair of spectral bins and can determine the projected uncertainty $\sigma\left[\Delta\theta\right]$ by combining measurements in multiple bins. Employing 70 spectral bins from the main diagonal allows to decrease the astrometric uncertainty by a factor of 8.3, while using extra bins from two adjacent diagonals, which are expected to have non-zero visibility, improved this value to 10.0. We consider this as an important step to demonstrate the advantages of frequency multiplexing and the viability of the approach. We also note that the number of spectral bins can be scaled up by a large factor employing larger and multiple sensors, and that it is not unusual to have tens of thousands of spectral channels on modern astronomical spectrographs [53].
Heisenberg-Gabor limit: It is instructive to compare the achieved resolution to the Heisenberg Uncertainty Principle limit or, equivalently, to the Heisenberg-Gabor limit. We note that the limits do not constrain the resolution for one of the observables alone, but restrict the product of the uncertainty of two observables, taking the form of:
$\Delta E\Delta t\geq\frac{\hbar}{2};~~~~\Delta f\Delta t\geq\frac{1}{4\pi},$ (4.2)
where $\hbar$ is Planck’s reduced constant, $\Delta E$, $\Delta f$, and $\Delta t$ are respectively the standard deviations for measurements of energy, frequency, and time. In our case of 640 nm wavelength and assuming 40 ps & 40 pm rms resolutions [44], $\Delta f\Delta t$ is a factor of 14.7 larger than the Heisenberg-Gabor limit, interpreting the spectral resolution uncertainty as frequency uncertainty. In general, this comparison is a useful benchmark for the indistinguishability of photons if they hit the same spectral bins in the two spectrometers. Experimentally, this means that the contrast of the corresponding HBT peak cannot be larger than $100\%/14.7=6.8\%$. The measured contrast of about 3-4% is indeed limited by the finite spectral and temporal resolutions of the spectrometer.
Entanglement swapping: Frequency multiplexing offers a powerful route to scale up entanglement swapping protocols in quantum communication, where SPDC sources are widely used to generate entangled photon pairs. A key limitation of SPDC is the generation of multi-pair events, which degrade quantum correlations; for QKD, the double-pair emission probability must typically remain below 5%, with even tighter constraints for other quantum applications [54, 55].
To enhance the rate of successful entanglement swaps, one promising strategy is to employ a broadband entanglement source, which can be viewed as comprising many narrowband emitters, multiplexed across frequency. The spectrometer described here can function as a multichannel entanglement-swapping platform, simultaneously receiving inputs from two entangled-photon sources. Each half of the sensor acts as an independent spectrometer, with pixels symmetrically paired across the array to monitor the same wavelength. Our configuration effectively enables up to 128 parallel Bell-state measurements (assuming 256 pixels in total are available) — corresponding to the number of spectral channels — thus increasing the entanglement swap rate by up to two orders of magnitude.
Given excellent timing resolution of the spectrometer, this architecture is particularly compatible with phase-encoded time-bin qubits [56], in which each qubit is encoded as a photon in a coherent superposition of early and late time bins. Entanglement swapping is then identified by two-photon coincidence events within the same spectral bin [57], which inform the remote parties of the spectral channel hosting the successfully entangled pair for subsequent quantum operations.
Future advancements in fast sensor technologies are expected to significantly enhance temporal resolution, potentially achieving resolutions as fine as 5 ps rms for SPAD devices, as recently demonstrated in [58]. Similarly, superconducting nanowire single-photon detectors (SNSPDs) are rapidly advancing toward temporal resolutions in the single-digit picosecond regime. However, these detectors currently require cryogenic temperatures, and their scalability to multi-channel and two-dimensional detector arrays has not yet been fully established [59, 60, 61].
Spectral resolution improvements down to a few picometers can be realized by employing an Echelle spectrometer configuration [62]. In this arrangement, high-order diffraction provides enhanced spectral resolution, while a first-order grating vertically separates overlapping spectral orders. The resulting output, a sequence of parallel spectral stripes, necessitates the use of a two-dimensional array of single-photon detectors to achieve high spectral resolution across an extensive wavelength range.
Collectively, these enhancements in temporal and spectral resolutions could restore visibility close to the theoretical limit, 100%, increasing it with respect to the present level by a factor of 20–40. This will translate to the same scale improvement of achievable astrometric resolution for the quantum-assisted astrometry applications and improved performance of the Bell-state measurements in quantum networking applications.
Ultimately, our realization of massively frequency-multiplexed two-photon correlation measurements presents a transformative advancement, with the potential to accelerate the development of quantum sensing and quantum communication technologies by orders of magnitude.
5 Conclusion
We have demonstrated the first observation of the HBT effect for broadband light dispersed in a fast spectrometer across a wide, continuous spectral range of 10 nm, utilizing a large number, of the order of 100, of spectral channels. After precise calibration, our spectrometer achieved exceptional temporal and spectral resolutions of 40 ps and 40 pm, respectively. This enabled the observation of thermal photon bunching with an average contrast of approximately 3%, consistent with theoretical expectations. The achieved temporal and spectral filtering approaches the Heisenberg–Gabor limit, making our spectrometer highly applicable to diverse fields such as quantum communications, quantum metrology, spectroscopy, and quantum-assisted telescopy.
Analyzing the spectrometer’s performance specifically in the context of phase-sensitive intensity interferometry, we found that combining about 200 spectral bin combinations exhibiting non-zero contrast could enhance astrometric precision by a factor of 10. Furthermore, we have outlined the significant potential gains achievable through entanglement swapping in quantum networks, where frequency multiplexing could dramatically enhance the throughput and robustness of entangled photon distribution.
Future advancements in temporal and spectral resolution promise further significant improvements, potentially enabling near-ideal visibility of two-photon interference phenomena. Ultimately, our realization of massively frequency-multiplexed two-photon correlation measurements presents a transformative advancement, with the potential to accelerate the development of quantum sensing and quantum communication technologies by orders of magnitude.
Acknowledgments
This research was supported by the Czech Science Foundation (GACR) under Project No. 25-15534M and the Grant Agency of the Czech Technical University in Prague, Grant No. SGS24/063/OHK4/1T/14. This work was also supported by the EPFL internal project “High-speed multimodal super-resolution microscopy with SPAD arrays” and DOE/LLNL project “The 3DQ Microscope”. The authors thank the Prague Observatory and Jakub Rozehnal for providing some of the optical equipment used in the experiments.